OpenStudy (anonymous):

1) A=3, B=2, C=1, H=8, Y=4 A*2/B-C*A*Y+18/C-3*H

OpenStudy (goformit100):

@Abbie23

OpenStudy (anonymous):

\[3*2/2 - 1*3*4 + 18/1 - 3*8\] can you do it now?

OpenStudy (anonymous):

but let me know how do you do it nphuongsun93

OpenStudy (anonymous):

substitute A,B,C,H,Y equals to the value you mentioned

OpenStudy (anonymous):

nphunongsun93 hello again, you know, i have done the practice at home before and I got a value, so I posted It here to see if your final result is the same that my final result, but well thanks for try to explain....

OpenStudy (anonymous):

\[\frac{2*3}{2}+\frac{18}{1}-(3*8)-(3*1*4)=-15 \]